How do I find the solutions when given:
y=x2+4.6+4.7
and y=3.3x-3.7

Answer by stanbon(75887)

How do I find the solutions when given:
y=x2+4.6+4.7
and y=3.3x-3.7
------
Substitute one y for the other:
x^2 + 4.6x + 4.7 = 3.3x - 3.7
------
x^2 + 1.3x + 8.4 = 0
-----
x = [-1.3 +- sqrt(1.69-4*8.4)]/2
---
x = [-1.3 +-sqrt(-31.91)]/2
----
That gives 2 complex number solution for "y".
That looks rather strange. Did you post the
problem correctly?
Cheers,
Stan H.
============